OpenStudy (anonymous):

Question 1.1. Solve, using the substitution method. j + k = 3 j – k = 7 The solution is (8, 1) The solution is (5, –2) There is no solution. There are an infinite number of solutions. Question 2.2. Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14 There is no solution. (12, 14) (14, 12) There are an infinite number of solutions. Question 3.3. Solve, using the substitution method. 3x + 2y = 8 y = x – 6 (0, –6) (3, –3) (4, –2) (14, 8)

OpenStudy (anonymous):

4. What is the x-coordinate of the solution to the system of equations? _________ Question 5. 5. What is the y-coordinate of the solution to the system of equations? _________

OpenStudy (anonymous):

1 question per post, so I'll just answer the first one.

OpenStudy (anonymous):

ok this is just some homework im having toruble with

OpenStudy (anonymous):

We can rearrange the first equation: j + k = 3 Subtract 'k' to both sides: j = -k + 3 Now we can plug this in for 'j' in the 2nd equation: j - k = 7 -k + 3 - k = 7 Simplify: -2k + 3 = 7 Subtract 3 to both sides, what's 7 - 3?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, so we have: -2k = 4 Now divide 4 / -2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yes, so k = -2 Now we plug this back into any of the two equations to solve for 'j': j + k = 3 j + -2 = 3 Add 2 to both sides, what's 3 + 2?

OpenStudy (anonymous):

You there? @Blackula123

OpenStudy (anonymous):

sorry my mom called me

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Yes, so 'j' is 5. And 'k' is -2. So we have (5, -2).

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

thankyou